A rigid vessel contains 6 kg of wet steam at 0.4 MPa with a quality of 75% until the steam


has a temperature of 300C. Determine the initial internal energy and the final pressure.

1
Expert's answer
2022-02-08T02:10:01-0500

at wet steam at 0.4 MPa with a quality of 75% 

 Final Temperature T2= 4600C from steam table

Internal Energy

"\delta E= m\times C_p dt =6\times 1\times (733-573) =960\frac{kJ}{kg}"

"P2 =P1\times 0.75 =0.4\times 0.75 = 0.3 Mpa"


as we konw,


"P_1T_1=P_2T_2"


"p_2= \frac{0.4\times573}{733}=0.3Mpa"